Consumer Banker Association released a report showing the lengths of automobile leases for new automobiles. The results are a(a) Use the midpoint of each class, and call the midpoint of the last class 66.5 months, for purposes of computing the expected lease term. Also find the standard deviation of the distribution. (Round your answers to two decimal places.)
Answer #1
x P(x) x* P(x) x square x 2* P(x)
18.5 0.121 2.24 342.25 41.41225
30.5 0.400 12.20 930.25 372.100
42.5 0.259 11.01 1806.25 467.819
54.5 0.215 11.72 2970.25 638.604
66.5 0.005 0.33 4422.25 22.111
sum 37.50 10471.25 1542.046

E(x) = sum of x. P(x)

Expected lease term =  2.24 + 12.20 + 11.01 + 11.72 + 0.33 = 37.50

Standard deviation

0 = 2 * p{) - 12 = Ꮩ1542.046 - 37,50

o 11.666

Standard deviation = 11.67